A Complete Bibliography of Publications in the International Journal of Applied Cryptography

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

02 January 2019 Version 1.02

Title word cross-reference

112 [BKK^{+}12]. π [FFGL10].

-bit [BKK^{+}12].

access [NNO^{+}17, SS09]. Achieving [CLL^{+}10]. active [BdMM08]. advanced [SS09, PB14]. adversaries [BdMM08]. adversary [PCR^{+}09]. AES [GSSC17, HSS09]. against [BM08, BdMM08, JLP10, Leu08]. agreement [MU09, MU10]. algebraic [HSS09]. ALPHA [HSS09]. among [HW08]. analysis [Cre10, HY08, LBM14, PR10, SS09, ACJT09]. Anonymous [BdMM08]. APOP [Leu08]. application [HSS09, JP13]. applications [SLdW12]. approach [LBM14]. arbitrary [BCS17]. Asiacrypt [ACJT09]. aspects [PR10]. assumption [CS13]. assumptions [KP17]. asymmetric [BCS17]. attack [Cre10, HN09, JLP10, LBM14, Leu08, PB14]. attacks [BM08, GSSC17, HS10]. Attribute [CG17, HJSNS12, AA14, EMO^{+}10, EMR12]. Attribute-based [CG17, HJSNS12, AA14, EMO^{+}10, EMR12]. auctions’ [DGK09]. authenticated [ACS17]. authentication [BdMM08, GMS09, Leu08, SS09]. authorisation [Tan12]. automatic [Cre10]. aware [GMS09].

backup [CML12]. backward [AA14]. based [AA14, BCS17, Boy08, CS13, CG17, EMO^{+}10, EMR12, GMS09, HJSNS12, ISW17, JP13, LBM14, Leu08, PSWH08, PR10, SOO10]. between [SEV17]. bilinear [CS13]. bit [BKK^{+}12]. blind [AO12]. block [YKP13, ZSWF10]. bounded [CG17]. browser [GMS09]. browser-based [GMS09].

call [HS10]. Camellia [YKP13]. case [GSSC17]. CCA [ISW17]. CCA-secure [ISW17]. CCA2 [BCS17]. certificateless [HW08]. chains [CLL^{+}10]. challenge [Leu08]. challenge-response [Leu08]. channel [PR10]. checking [FYL17]. Chosen [SLdW12]. Chosen-prefix [SLdW12]. cipher [HN09, YKP13, ZSWF10]. ciphertext [CML12, EMO^{+}10]. ciphertext-policy [EMO^{+}10]. ciphertexts [ACS17, KP17]. cloning [HJSNS12]. collisions [SLdW12]. commit [Lip17]. commit-and-prove [Lip17]. committing [AO12]. compared [Boy08]. Comparing [MU09]. comparison [DGK08, DGK09, Veu14]. composability [OMO08]. composable [AO12]. computation [JP13]. computing [JP13]. confirmation [GMS10]. consoles [BKK^{+}12]. constant [BdMM08, EMO^{+}10]. constant-cost [BdMM08]. constrained [ACS17]. content [SS09]. correction [DGK09]. corruptions [BM08]. cost [BdMM08]. critical [SS09]. cryptanalysis [ZSWF10]. cryptography [HW08]. curve [BKK^{+}12]. curves [CS12, Far14].

data [FYL17]. databases [YBDD09]. datasets [YWPZ09]. decisional [CS13]. decryption [HKK^{+}10, ISW17]. Defending [JLP10]. Delayed [JM08]. derivation [CS12]. Designing [HS10]. devices [ACS17]. different [Tan12]. Diffie [CS13, CS12, MU10]. disclosure [JM08]. discrete [BKK^{+}12]. disjointness [YWPZ09]. distance [JP13]. distinguishing [HN09]. division [Veu14]. document [JP13]. Double [CML12]. drive [SS09]. drive-host [SS09]. Dynamic [AA14, EMR12].

eCKs [Cre10]. efficiency [CLL^{+}10]. Efficient [DSMRY12, PSWH08, HLW08, Lip17, DGK09]. elliptic [BKK^{+}12, CS12]. Encrypted [Veu14, YBDD09]. Encryption [PB14, ACS17, BCS17, Boy08, CS13, CG17, DGK08, EMO^{+}10, HJSNS12, ISW17, Tan12]. end [CS12]. enhanced [NNO^{+}17]. ephemeral [MU10]. EphemeralKeyReveal [Cre10]. equality [Tan12]. evolution [JLP10]. exchange [BCNP09, BM08, OMO08]. explicit [GMS10]. exposure [ISW17]. extended [BCS17]. extensions [KP17]. extraction [CS12].

Fair [HKK^{+}10, OMO08]. fast [FFGL10]. fault [GSSC17]. feasibility [PCRS10]. FHMQV [SEV17]. five [HSS09]. five-round [HSS09]. forest [PB14]. framework [AO12, FFGL10, HY08, OMO08]. frameworks [Boy08]. fully [CG17]. function [KP17]. function-private [KP17]. functionalities [JP13]. functions [FFGL10].

game [BKK^{+}12]. games [HY08]. granularity [Tan12]. group [AA14, ACJT09, BM08].

Hamming [JP13]. handling [ACS17]. hash [FFGL10]. Hashing [Far14]. Hellman [CS13, CS12, MU10]. Hessian [Far14]. hierarchical [CS13]. high [CLL^{+}10]. HMQV [SEV17]. Homomorphic [DGK08, CG17]. host [SS09]. hybrid [CS13].

IBE [KP17]. ID [SOO10]. ID-based [SOO10]. ideal [BNV17]. identification [KH08]. identity [Boy08, CS13, ISW17]. identity-based [Boy08, CS13, ISW17]. improvement [SS09]. information [PR10]. information-based [PR10]. inputs [CG17]. integer [Veu14]. integers [LN14]. integrity [FYL17]. intersection [DSMRY12].

key [BCNP09, BM08, BdMM08, CS12, HJSNS12, ISW17, Leu08, MU09, MU10, Tan12]. key-lookup [BdMM08]. key-recovery [Leu08]. keys [KP17, MU10]. knowledge [Lip17].

lattice [PSWH08]. lattice-based [PSWH08]. lattices [BNV17]. layered [HY08]. learning [LBM14]. length [BCS17, EMO^{+}10]. line [DGK09]. linear [HN09, KP17]. local [AA14]. log [NNO^{+}17]. logarithm [BKK^{+}12]. long [ACS17]. lookup [BdMM08].

MAC [HSS09]. machine [LBM14]. management [NNO^{+}17]. MD5 [Leu08, SLdW12]. MD5-based [Leu08]. mechanism [NNO^{+}17]. memory [ACS17]. message [BCS17, GMS10, PCR^{+}09, PCRS10]. mobile [PCR^{+}09, WOS09]. mode [CML12]. model [AA14, BCNP09]. models [HW08, MU09]. move [HLW08]. MR2444651 [ACJT09]. multidimensional [HN09]. Multiuser [YBDD09]. mutual [GMS09, PR10].

NAXOS [Cre10]. networks [PCRS10]. nominative [HLW08, NNO^{+}17]. non [AO12]. non-committing [AO12]. Notions [LN14]. number [CG17].

offline [WOS09]. on-line [DGK09]. One [BCNP09, ACJT09, BDE^{+}13, HLW08]. one-move [HLW08]. One-round [BCNP09]. one-time [BDE^{+}13]. optimality [PCRS10]. optimistic [OMO08]. oracles [EMR12, SOO10].

p [YKP13]. p-Camellia [YKP13]. p-SMS4 [YKP13]. Parallelisable [YKP13]. parties [HKK^{+}10]. password [JM08]. peer [MU09]. Perfectly [PCR^{+}09]. perspective [BNV17]. pirate [JLP10]. platforms [WOS09]. policy [EMO^{+}10]. popular [ACJT09]. possibility [PCRS10]. post [MU09]. post-specified [MU09]. Power [LBM14, KH08]. Practical [CS13, Leu08, BNV17, Boy08, PR10]. pre [MU09]. pre- [MU09]. Preface [BW10]. prefix [SLdW12]. Preventing [GSSC17]. prime [BKK^{+}12]. privacy [FYL17, NNO^{+}17]. privacy-enhanced [NNO^{+}17]. private [DSMRY12, KP17, YBDD09, YWPZ09]. problem [BKK^{+}12]. profiling [PB14]. property [HSS09]. proposal [CML12]. protocol [CS12, Cre10, OMO08]. protocols [GMS09, GMS10, HY08, MU10, SEV17]. provably [GMS09]. prove [Lip17]. Prover [Lip17]. Prover-efficient [Lip17]. Public [Tan12, FYL17, LW17].

queries [YBDD09].

Random [PB14, EMR12, SOO10]. randomisation [GSSC17]. Randomness [CS12, LW17]. recognition [GMS10]. recoverability [GMS10]. recovery [Leu08]. reduction [BKK^{+}12]. registration [BM08]. relation [HW08]. reliable [PCR^{+}09, PCRS10]. Remarks [ACJT09]. remote [FYL17]. resistance [ISW17]. response [Leu08]. results [ZSWF10]. reusing [MU10]. revocable [ISW17]. revocation [AA14]. RFID [BdMM08, CLL^{+}10]. RFID-tagged [CLL^{+}10]. robust [DSMRY12]. round [BCNP09, HSS09]. RSA [LN14]. scheme [ACJT09, BDE^{+}13, EMO^{+}10, FYL17, HLW08, PSWH08]. schemes [CS13, ISW17, KH08, Tan12]. secret [CS12]. Secure [JP13, WOS09, BCS17, CML12, DGK08, DGK09, FFGL10, GMS09, ISW17, PCR^{+}09, PCRS10, Veu14, YWPZ09]. Securing [BM08]. security [BDE^{+}13, CLL^{+}10, HY08, HW08, OMO08]. semi [HKK^{+}10]. semi-trusted [HKK^{+}10]. separation [SEV17]. Session [Cre10]. Session-StateReveal [Cre10]. set [DSMRY12, HS10]. set-up [HS10]. Shannon [HN09]. shorter [KP17]. shortest [BNV17]. side [PR10]. Sieving [BNV17]. signature [AA14, ACJT09, BDE^{+}13, HLW08, PSWH08]. signatures [AO12, NNO^{+}17, SOO10]. signcryption [EMR12]. similarity [JP13]. SIP [HS10]. sloppy [BKK^{+}12]. sloth [LW17]. SMS4 [YKP13, ZSWF10]. SNARKs [Lip17]. Solving [BKK^{+}12]. Some [ZSWF10]. specifications [HY08]. specified [MU09]. Sponge [BCS17]. Sponge-based [BCS17]. SSO [NNO^{+}17]. standard [AA14, BCNP09, PB14]. StateReveal [Cre10]. strong [BM08]. stronger [Cre10]. Strongly [SOO10]. study [GSSC17]. superdistribution [WOS09]. supply [CLL^{+}10]. supporting [BdMM08, Tan12]. synchronous [PCRS10]. system [SS09]. systems [NNO^{+}17].

tagged [CLL^{+}10]. tapestry [Boy08]. technique [ACS17]. test [Tan12]. tests [YWPZ09]. Theoretical [PR10]. third [HKK^{+}10]. threshold [HKK^{+}10]. time [BDE^{+}13]. tolerating [PCR^{+}09]. transmission [PCR^{+}09, PCRS10]. trusted [HKK^{+}10]. Trustworthy [LW17]. trx [LW17]. TWISTER [FFGL10].

Unconditionally [PCRS10, YWPZ09]. undirected [PCRS10]. unforgeable [SOO10]. unicorn [LW17]. universal [OMO08]. universally [AO12]. unlinkability [AA14]. User [GMS09, FYL17]. User-aware [GMS09]. using [BKK^{+}12, Cre10, GSSC17].

variants [YKP13]. various [HW08]. vectors [BNV17]. verifier [AA14]. verifier-local [AA14]. version [BCS17].

Winternitz [BDE^{+}13]. without [EMR12, HJSNS12, SOO10].

zero [Lip17]. zero-knowledge [Lip17].

References

Ali:2014:DAB

[AA14] Syed Taqi Ali and B. B. Amberker. Dynamic attribute-based group signature with verifierlocal revocation and backward unlinkability in the standard model. International Journal of Applied Cryptography. IJACT, 3 (2):148–165, 2014. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Ateniese:2009:RAO

[ACJT09] Giuseppe Ateniese, Jan Camenisch, Marc Joye, and Gene Tsudik. Remarks on “Analysis of one popular group signature scheme” in Asiacrypt 2006 [MR2444651]. International Journal of Applied Cryptography. IJACT, 1(4):320–322, 2009. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Agrawal:2017:NAE

[ACS17] Megha Agrawal, Donghoon Chang, and Somitra Kumar Sanadhya. A new authenticated encryption technique for handling long ciphertexts in memory constrained devices. International Journal of Applied Cryptography. IJACT, 3(3):236–261, 2017. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic). URL https://

www.inderscienceonline.com/
/doi/pdf/10.1504/IJACT.2017.
086223.

Abe:2012:FUC

[AO12] Masayuki Abe and Miyako Ohkubo. A framework for universally composable noncommitting blind signatures. International Journal of Applied Cryptography. IJACT, 2(3):229– 249, 2012. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Boyd:2009:ORK

[BCNP09] Colin Boyd, Yvonne Cliff, Juan M. González Nieto, and Kenneth G. Paterson. One-round key exchange in the standard model. International Journal of Applied Cryptography. IJACT, 1 (3):181–199, 2009. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Bansal:2017:SBC

[BCS17] Tarun Kumar Bansal, Donghoon Chang, and Somitra Kumar Sanadhya. Sponge-based CCA2 secure asymmetric encryption for arbitrary length message (extended version). International Journal of Applied Cryptography. IJACT, 3(3):262–287, 2017. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic). URL https://

                                        www.inderscienceonline.com/
                                        /doi/pdf/10.1504/IJACT.2017.
                                        086222.

Buchmann:2013:SWO

[BDE^{+}13] Johannes Buchmann, Erik Dahmen, Sarah Ereth, Andreas Hülsing, and Markus Rückert. On the security of the Winternitz one-time signature scheme. International Journal of Applied Cryptography. IJACT, 3 (1):84–96, 2013. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Burmester:2008:ARA

[BdMM08] M. Burmester, B. de Medeiros, and R. Motta. Anonymous RFID authentication supporting constant-cost key-lookup against active adversaries. International Journal of Applied Cryptography. IJACT, 1(2):79–90, 2008. CO-DEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Bos:2012:SBP

[BKK^{+}12] Joppe W. Bos, Marcelo E. Kaihara, Thorsten Kleinjung, Arjen K. Lenstra, and Peter L. Montgomery. Solving a 112-bit prime elliptic curve discrete logarithm problem on game consoles using sloppy reduction. International Journal of Applied Cryptography. IJACT, 2(3):212–228, 2012. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Bresson:2008:SGK

[BM08] Emmanuel Bresson and Mark Manulis. Securing group key exchange against strong corruptions and key registration attacks. International Journal of Applied Cryptography. IJACT, 1 (2):91–107, 2008. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Bos:2017:SSV

[BNV17] Joppe W. Bos, Michael Naehrig, and Joop Van De Pol. Sieving for shortest vectors in ideal lattices: a practical perspective. International Journal of Applied Cryptography. IJACT, 3 (4):313–329, 2017. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic). URL https://

www.inderscienceonline.com/
/doi/pdf/10.1504/IJACT.2017.
089353.

Boyen:2008:TIB

[Boy08] Xavier Boyen. A tapestry of identity-based encryption: practical frameworks compared. International Journal of Applied Cryptography. IJACT, 1(1):3–21, 2008. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Bao:2010:P

[BW10] Feng Bao and Guilin Wang. Preface. International Journal of Applied Cryptography. IJACT, 2 (1):1–2, 2010. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Clear:2017:ABF

[CG17] Michael Clear and Ciarán Mc Goldrick. Attribute-based fully homomorphic encryption with a bounded number of inputs. International Journal of Applied Cryptography. IJACT, 3(4): 363–376, 2017. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic). URL https://

www.inderscienceonline.com/
/doi/pdf/10.1504/IJACT.2017.
089356.

Cai:2010:AHS

[CLL^{+}10] Shaoying Cai, Yingjiu Li, Tieyan Li, Robert H. Deng, and Haixia Yao. Achieving high security and efficiency in RFID-tagged supply chains. International Journal of Applied Cryptography. IJACT, 2 (1):3–12, 2010. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Chakraborty:2012:DCM

[CML12] Debrup Chakraborty and Cuauhtemoc Mancillas-López. Double ciphertext mode: a proposal for secure backup. International Journal of Applied Cryptography. IJACT, 2(3):271–287, 2012. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Cremers:2010:SSS

[Cre10] Cas J. F. Cremers. Session-StateReveal is stronger than eCKs EphemeralKeyReveal: using automatic analysis to attack the NAXOS protocol. International Journal of Applied Cryptography. IJACT, 2(2):83–99, 2010. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Ciss:2012:REE

[CS12] Abdoul Aziz Ciss and Djiby Sow.

Randomness extraction in elliptic curves and secret key derivation at the end of Diffie–Hellman protocol. International Journal of Applied Cryptography. IJACT, 2(4):360–365, 2012. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Chatterjee:2013:PHH

[CS13] Sanjit Chatterjee and Palash Sarkar. Practical hybrid (hierarchical) identity-based encryption schemes based on the decisional bilinear Diffie–Hellman assumption. International Journal of Applied Cryptography. IJACT, 3(1):47–83, 2013. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Damgaard:2008:HES

[DGK08] Ivan Damgård, Martin Geisler, and Mikkel Krøigård. Homomorphic encryption and secure comparison. International Journal of Applied Cryptography. IJACT, 1 (1):22–31, 2008. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Damgaard:2009:CES

[DGK09] Ivan Damgård, Martin Geisler, and Mikkel Krøigård. A correction to ‘Efficient and secure comparison for on-line auctions’. International Journal of Applied Cryptography. IJACT, 1 (4):323–324, 2009. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Dachman-Soled:2012:ERP

[DSMRY12] Dana Dachman-Soled, Tal Malkin, Mariana Raykova, and Moti Yung. Efficient robust private set intersection. International Journal of Applied Cryptography. IJACT, 2(4):289–303, 2012. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Emura:2010:CPA

[EMO^{+}10] Keita Emura, Atsuko Miyaji, Kazumasa Omote, Akito Nomura, and Masakazu Soshi. A ciphertext-policy attribute-based encryption scheme with constant ciphertext length. International Journal of Applied Cryptography. IJACT, 2(1):46–59, 2010. CO-DEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Emura:2012:DAB

[EMR12] Keita Emura, Atsuko Miyaji, and Mohammad Shahriar Rahman. Dynamic attribute-based signcryption without random oracles. International Journal of Applied Cryptography. IJACT, 2 (3):199–211, 2012. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Farashahi:2014:HHC

[Far14] Reza Rezaeian Farashahi. Hashing into Hessian curves. International Journal of Applied Cryptography. IJACT, 3(2):139–147, 2014. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Fleischmann:2010:TFS

[FFGL10] Ewan Fleischmann, Christian Forler, Michael Gorski, and Stefan Lucks. TWISTERπ — a framework for secure and fast hash functions. International Journal of Applied Cryptography. IJACT, 2(1):68–81, 2010. CO-DEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Feng:2017:NPR

[FYL17] Yiteng Feng, Guomin Yang, and Joseph K. Liu. A new public remote integrity checking scheme with user and data privacy. International Journal of Applied Cryptography. IJACT, 3 (3):196–209, 2017. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic). URL https://

www.inderscienceonline.com/
/doi/pdf/10.1504/IJACT.2017.
086232.

Gajek:2009:UAP

[GMS09] Sebastian Gajek, Mark Manulis, and Jörg Schwenk. User-aware provably secure protocols for browser-based mutual authentication. International Journal of Applied Cryptography. IJACT, 1 (4):290–308, 2009. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Goldberg:2010:MRP

[GMS10] Ian Goldberg, Atefeh Mashatan, and Douglas R. Stinson. On message recognition protocols: recoverability and explicit confirmation. International Journal of Applied Cryptography. IJACT, 2 (2):100–120, 2010. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Ghosh:2017:PFA

[GSSC17] Shamit Ghosh, Dhiman Saha, Abhrajit Sengupta, and Dipanwita Roy Chowdhury. Preventing fault attacks using fault randomisation with a case study on AES. International Journal of Applied Cryptography. IJACT, 3 (3):225–235, 2017. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic). URL https://

www.inderscienceonline.com/
/doi/pdf/10.1504/IJACT.2017.
086231.

Hinek:2012:ABE

[HJSNS12] M. Jason Hinek, Shaoquan Jiang, Reihaneh Safavi-Naini, and Siamak F. Shahandashti. Attribute-based encryption without key cloning. International Journal of Applied Cryptography. IJACT, 2(3):250–270, 2012. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Hong:2010:FTD

[HKK^{+}10] Jeongdae Hong, Jinil Kim, Jihye Kim, Matthew K. Franklin, and Kunsoo Park. Fair threshold decryption with semi-trusted third parties. International Journal of Applied Cryptography. IJACT, 2 (2):139–153, 2010. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Huang:2008:EOM

[HLW08] Qiong Huang, Dennis Y. W. Liu, and Duncan S. Wong. An efficient one-move nominative signature scheme. International Journal of Applied Cryptography. IJACT, 1(2):133–143, 2008. CO-DEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Hakala:2009:MLD

[HN09] Risto M. Hakala and Kaisa Nyberg. A multidimensional linear distinguishing attack on the Shannon cipher. International Journal of Applied Cryptography. IJACT, 1(3):161–168, 2009. CO-DEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Hagalisletto:2010:DAS

[HS10] Anders Moen Hagalisletto and Lars Strand. Designing attacks on SIP call set-up. International Journal of Applied Cryptography. IJACT, 2(1):13–22, 2010. CO-DEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Huang:2009:FRA

[HSS09] Jianyong Huang, Jennifer Seberry, and Willy Susilo. A five-round algebraic property of AES and its application to the ALPHA–MAC. International Journal of Applied Cryptography. IJACT, 1(4):264–289, 2009. CO-DEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Huang:2008:RAV

[HW08] Qiong Huang and Duncan S.

Wong. On the relation among various security models for certificateless cryptography. International Journal of Applied Cryptography. IJACT, 1(2):108– 119, 2008. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Herzberg:2008:LGF

[HY08] Amir Herzberg and Igal Yoffe. The layered games framework for specifications and analysis of security protocols. International Journal of Applied Cryptography. IJACT, 1(2):144–159, 2008. CO-DEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Ishida:2017:CSR

[ISW17] Yuu Ishida, Junji Shikata, and Yohei Watanabe. CCA-secure revocable identity-based encryption schemes with decryption key exposure resistance. International Journal of Applied Cryptography. IJACT, 3(3):288–311, 2017. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic). URL https://

www.inderscienceonline.com/
/doi/pdf/10.1504/IJACT.2017.
086229.

Jin:2010:DAP

[JLP10] Hongxia Jin, Jeffrey Lotspiech, and Serdar Pehlivanoglu. Defending against the pirate evolution attack. International Journal of Applied Cryptography. IJACT, 2(1):23–34, 2010. CO-DEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Jakobsson:2008:DPD

[JM08] Markus Jakobsson and Steven Myers. Delayed password disclosure. International Journal of Applied Cryptography. IJACT, 1 (1):47–59, 2008. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Jarrous:2013:SCF

[JP13] Ayman Jarrous and Benny Pinkas. Secure computation of functionalities based on Hamming distance and its application to computing document similarity. International Journal of Applied Cryptography. IJACT, 3 (1):21–46, 2013. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Kurosawa:2008:PIS

[KH08] Kaoru Kurosawa and Swee-Huay Heng. The power of identification schemes. International Journal of Applied Cryptography. IJACT, 1(1):60–69, 2008. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Kurosawa:2017:IFP

[KP17] Kaoru Kurosawa and Le Trieu Phong. IBE and functionprivate IBE under linear assumptions with shorter ciphertexts and private keys, and extensions. International Journal of Applied Cryptography. IJACT, 3 (3):210–224, 2017. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic). URL https://

www.inderscienceonline.com/

                                        /doi/pdf/10.1504/IJACT.2017.
                                        086224.

Lerman:2014:PAA

[LBM14] Liran Lerman, Gianluca Bontempi, and Olivier Markowitch. Power analysis attack: an approach based on machine learning. International Journal of Applied Cryptography. IJACT, 3 (2):97–115, 2014. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Leurent:2008:PKR

[Leu08]  Gaëtan Leurent.  Practi-

cal key-recovery attack against APOP, an MD5-based challengeresponse authentication. International Journal of Applied Cryptography. IJACT, 1(1):32– 46, 2008. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Lipmaa:2017:PEC

[Lip17] Helger Lipmaa. Prover-efficient commit-and-prove zero-knowledge SNARKs. International Journal of Applied Cryptography. IJACT, 3(4):344–362, 2017. CO-DEN ???? ISSN 1753-0563 (print), 1753-0571 (elec-

                                        tronic). URL https://www.
                                        inderscienceonline.com//doi/
                                        pdf/10.1504/IJACT.2017.089355.

Loebenberger:2014:NRI

[LN14] Daniel Loebenberger and Michael Nüsken. Notions for RSA integers. International Journal of Applied Cryptography. IJACT, 3 (2):116–138, 2014. CODEN ????

ISSN 1753-0563 (print), 1753-0571 (electronic).

Lenstra:2017:TPR

[LW17] Arjen K. Lenstra and Benjamin Wesolowski. Trustworthy public randomness with sloth, unicorn, and trx. International Journal of Applied Cryptography. IJACT, 3 (4):330–343, 2017. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic). URL https://

www.inderscienceonline.com/
/doi/pdf/10.1504/IJACT.2017.
089354.

Menezes:2009:CPP

[MU09] Alfred Menezes and Berkant Ustaoglu. Comparing the preand post-specified peer models for key agreement. International Journal of Applied Cryptography. IJACT, 1(3):236–250, 2009. CO-DEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Menezes:2010:REK

[MU10] Alfred Menezes and Berkant Ustaoglu. On reusing ephemeral keys in Diffie–Hellman key agreement protocols. International Journal of Applied Cryptography. IJACT, 2(2):154–158, 2010. CO-DEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Nakagawa:2017:PEA

[NNO^{+}17] Sanami Nakagawa, Takashi Nishide, Eiji Okamoto, Keita Emura, Goichiro Hanaoka, Yusuke Sakai, and Akihisa Kodate. A privacy-enhanced access log management mechanism in SSO systems from nominative signatures. International Journal of Applied Cryptography. IJACT, 3 (4):394–406, 2017. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic). URL https://

www.inderscienceonline.com/
/doi/pdf/10.1504/IJACT.2017.
089373.

Okada:2008:OFE

[OMO08] Yusuke Okada, Yoshifumi Manabe, and Tatsuaki Okamoto. An optimistic fair exchange protocol and its security in the universal composability framework. International Journal of Applied Cryptography. IJACT, 1 (1):70–77, 2008. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Patel:2014:RFP

[PB14] Hiren Patel and Rusty O. Baldwin. Random forest profiling attack on Advanced Encryption Standard. International Journal of Applied Cryptography. IJACT, 3(2):181–194, 2014. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Patra:2009:PRS

[PCR^{+}09] Arpita Patra, Ashish Choudhary, C. Pandu Rangan, Kannan Srinathan, and Prasad Raghavendra. Perfectly reliable and secure message transmission tolerating mobile adversary. International Journal of Applied Cryptography. IJACT, 1(3):200–224, 2009. CO-DEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Patra:2010:URS

[PCRS10] Arpita Patra, Ashish Choudhury, C. Pandu Rangan, and Kannan Srinathan. Unconditionally reliable and secure message transmission in undirected synchronous networks: possibility, feasibility and optimality. International Journal of Applied Cryptography. IJACT, 2(2): 159–197, 2010. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Prouff:2010:TPA

[PR10] E. Prouff and M. Rivain. Theoretical and practical aspects of mutual information-based side channel analysis. International Journal of Applied Cryptography. IJACT, 2(2):121–138, 2010. CO-DEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Plantard:2008:ELB

[PSWH08] Thomas Plantard, Willy Susilo, Khin Than Win, and Qiong Huang. Efficient lattice-based signature scheme. International Journal of Applied Cryptography. IJACT, 1(2):120–132, 2008. CO-DEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Sarr:2017:SBF

[SEV17] Augustin P. Sarr and Philippe Elbaz-Vincent. On the separation between the FHMQV and HMQV protocols. International Journal of Applied Cryptography. IJACT, 3(4):377–393, 2017. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic). URL https://

                                        www.inderscienceonline.com/
                                        /doi/pdf/10.1504/IJACT.2017.
                                        089357.

Stevens:2012:CPC

[SLdW12] Marc Stevens, Arjen K. Lenstra, and Benne de Weger. Chosenprefix collisions for MD5 and applications. International Journal of Applied Cryptography. IJACT, 2(4):322–359, 2012. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Sato:2010:SUI

[SOO10] Chifumi Sato, Takeshi Okamoto, and Eiji Okamoto. Strongly unforgeable ID-based signatures without random oracles. International Journal of Applied Cryptography. IJACT, 2(1):35–45, 2010. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Sui:2009:CAI

[SS09] Jiayuan Sui and Douglas R. Stinson. A critical analysis and improvement of advanced access content system drive-host authentication. International Journal of Applied Cryptography. IJACT, 1(3):169–180, 2009. CO-DEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Tang:2012:PKE

[Tan12] Qiang Tang. Public key encryption schemes supporting equality test with authorisation of different granularity. International Journal of Applied Cryptography. IJACT, 2(4):304–321, 2012. CO-DEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Veugen:2014:EID

[Veu14] Thijs Veugen. Encrypted integer division and secure comparison. International Journal of Applied Cryptography. IJACT, 3 (2):166–180, 2014. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Wolf:2009:SOS

[WOS09] Marko Wolf, André Osterhues, and Christian Stüble. Secure offline superdistribution for mobile platforms. International Journal of Applied Cryptography. IJACT, 1(4):251–263, 2009. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Yang:2009:MPQ

[YBDD09] Yanjiang Yang, Feng Bao, Xuhua Ding, and Robert H. Deng. Multiuser private queries over encrypted databases. International Journal of Applied Cryptography. IJACT, 1(4):309–319, 2009. CO-DEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Yap:2013:PVC

[YKP13] Huihui Yap, Khoongming Khoo, and Axel Poschmann. Parallelisable variants of Camellia and SMS4 block cipher: p-Camellia and p-SMS4. International Journal of Applied Cryptography. IJACT, 3(1):1–20, 2013. CO-DEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Ye:2009:USD

[YWPZ09] Qingsong Ye, Huaxiong Wang, Josef Pieprzyk, and Xian-Mo Zhang. Unconditionally secure disjointness tests for private datasets. International Journal of Applied Cryptography. IJACT, 1(3):225–235, 2009. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Zhang:2010:SRC

[ZSWF10] Wentao Zhang, Bozhan Su, Wenling Wu, and Dengguo Feng. Some results on cryptanalysis of SMS4 block cipher. International Journal of Applied Cryptography. IJACT, 2(1):60–67, 2010. CO-DEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).